# Characterization of an MPPC-Based Scintillator Telescope and Measurement of Cosmic Muon Angular Distribution

### v1

@@ V-A The Angular Exponent $n$
This measured value lies reasonably close to the commonly used theoretical approximation of $n=2$ [undefi], which is often employed as a first-order description of the angular distribution. Deviations from this idealized value are expected in real measurements due to factors such as atmospheric attenuation, detector geometry, and local environmental conditions.
@@
As shown in the table, our result of $n=1.44\pm 0.06$ falls within the general range of values reported in the literature, though it is notably lower than the theoretical approximation of $n=2$. The agreement with the study by Venterea and Ekka ($1.39\pm 0.01$) is particularly strong. This consistency suggests that our detector geometry and local conditions are comparable to those in their experiment, providing high confidence in the validity of our experimental method and analysis.
@@ V-B Analysis of the Isotropic Background Component (B or d parameter)
This constant background rate aligns well with the general expectation that at large zenith angles, as the true muon flux significantly diminishes, ambient radiation and other local effects become the dominant signal components [undefb, undefk]. This background is likely composed of two main sources: a small, residual rate of accidental coincidences from detector dark counts, and a physical component from cosmic ray side-showers or local ambient radiation that can trigger the detectors from horizontal directions. The inclusion of this $d$ parameter was essential for achieving a good statistical fit, as demonstrated by the significantly lower Reduced Chi-Squared value of the Schwerdt model ($\chi^{2}_{\nu}=0.88$) compared to the simple $\cos^{2}(\theta)$ model ($\chi^{2}_{\nu}=8.36$).
@@ VI Conclusion
The primary goal of the experiment was to measure the angular distribution of the muon flux. The measured rate was found to decrease with increasing zenith angle, consistent with theory. A comparative analysis of four prominent models from the literature was performed to quantitatively describe this distribution. While the Schwerdt model provided the lowest chi-squared value, the General $\cos^{n}(\theta)$ model was utilized to determine the physical angular dependence. From this fit, the angular exponent of the cosmic ray muon flux was measured to be $n=1.44\pm 0.06$. This result is in excellent agreement with values reported in the published literature (specifically Venterea and Ekka [undefb]), though it deviates from the standard theoretical approximation of $n=2$ due to the detector’s geometric acceptance and local environmental factors.
@@ Appendix A Experimental Operating Parameters
*Note: The threshold was set approximately $100~\mu\text{V}$ below the mean 3 p.e. amplitude to maximize detection efficiency while rejecting 2 p.e. noise.

### v2

@@ V-A The Angular Exponent $n$
This measured value lies reasonably close to the commonly used theoretical approximation of $n=2$ [undefj], which is often employed as a first-order description of the angular distribution. Deviations from this idealized value are expected in real measurements due to factors such as atmospheric attenuation, detector geometry, and local environmental conditions.
@@
As shown in the Table IX, our result of $n=1.44\pm 0.06$ falls within the general range of values reported in the literature, though it is notably lower than the theoretical approximation of $n=2$. The agreement with the study by Venterea and Ekka ($1.39\pm 0.01$)  [undefb] is particularly strong. This consistency suggests that our detector geometry and local conditions are comparable to those in their experiment, providing high confidence in the validity of our experimental method and analysis.
@@ V-B Analysis of the Isotropic Background Component (B or d parameter)
This constant background rate aligns well with the general expectation that at large zenith angles, as the true muon flux significantly diminishes, ambient radiation and other local effects become the dominant signal components [undefb, undefl]. This background is likely composed of two main sources: a small, residual rate of accidental coincidences from detector dark counts, and a physical component from cosmic ray side-showers or local ambient radiation that can trigger the detectors from horizontal directions. The inclusion of this $d$ parameter was essential for achieving a good statistical fit, as demonstrated by the significantly lower Reduced Chi-Squared value of the Schwerdt model ($\chi^{2}_{\nu}=0.88$) compared to the simple $\cos^{2}(\theta)$ model ($\chi^{2}_{\nu}=8.36$).
@@ VI Conclusion
The primary goal of the experiment was to measure the angular distribution of the muon flux. The measured rate was found to decrease with increasing zenith angle, consistent with theory. A comparative analysis of four prominent models from the literature was performed to quantitatively describe this distribution. While the Schwerdt model provided the lowest chi-squared value, the General $\cos^{n}(\theta)$ model was utilized to determine the physical angular dependence. From this fit, the angular exponent of the cosmic ray muon flux was measured to be $n=1.44\pm 0.06$. This result is in excellent agreement with values reported in the published literature (specifically Venterea and Ekka), though it deviates from the standard theoretical approximation of $n=2$ due to the detector’s geometric acceptance and local environmental factors.
@@ Appendix A Experimental Operating Parameters
Note: The threshold was set approximately $100~\mu\text{V}$ below the mean 3 p.e. amplitude to maximize detection efficiency while rejecting 2 p.e. noise.